# A longitudinal Bayesian framework for estimating causal dose-response relationships

### v1

@@ 3.1 Example 1
$D$ is a continuous treatment variable, and $X_{1}$, $X_{2}$ are time-varying confounders and $U$ is a time-invariant confounder for $D$ with a nonlinear relationship in $Y$. The goal of this causal setting is to estimate the APO, i.e., the outcome at a specific dose level. We consider estimating the dose-response relationship based on a GEE model for GPS with the inclusion of covariates $X_{1it}$, $X_{2it}$ only. For the COV method, we fit a GEE model for the outcome of $\log Y_{i}$ with the inclusion of covariates $D_{it}$, the three-order spline of estimated GPS. As for WOR method, we fit the outcome mean with the three-order spline of $D_{1it}$ with the GPS as the weight. In the DP method, the new data are generated from the old dataset or new data (unseen) from the base measure, $G_{0}$. We generate these new data through residuals arising from the outcome model. In this case, we first draw each pair of $\left\{y^{s^{\prime}}_{ik},({x}^{s}_{ik},d^{s}_{ik})\right\}$, $k=1,\ldots,K_{i},i=1\ldots,N$, from the empirical distribution as the DP with $\alpha=0$, and then obtain the GPS, $e\left({x}^{s}_{ik};\widehat{\gamma}^{s}\right)$ from a mixed effect model with the inclusion of covariates $x^{s}_{1it}$, $x^{s}_{2it}$ to account for the effect of the unmeasured confounder (Graham et al., 2014), refitted to the newly sampled $\left\{({x}^{s}_{i},d^{s}_{i})\right\}$ data set. Then we simulate $y_{i}^{s}=\left\{y_{ik}^{s}\right\}_{k=1}^{K}$ from a DP model with the conditional base measure $G_{0}\equiv\mathcal{N}_{K}\left(\mu({x}^{s}_{i},e\left({x}^{s}_{i};\widehat{\gamma}^{s}\right)),1\right)$, where $\mu({x}^{s}_{i},e\left({x}^{s}_{i};\widehat{\gamma}^{s}\right))$ is the fitted value from a GEE model with $\log y^{s^{\prime}}_{i}$ with the inclusion of covariates $d^{s}_{i}$, the three-order spline of estimated GPS, $e\left({x}^{s}_{i};\widehat{\gamma}^{s}\right)$. In the stick breaking construction, we truncate the sample size at $J=500$. The weight is less than $10^{-4}$ after 500 iterations and can be deemed as negligible.
@@ 3.2 Example 2
Similar to the first example, we fit a Poisson GEE model for the outcome of $Y_{i}$ with the inclusion of covariates $D_{1it}$, the three-order spline of estimated GPS for COV method, and as for WOR method, we fit the outcome mean with the three-order spline of $D_{1it}$ with the GPS as the weight. For the DP approach, we use the same steps to sample the data with size 500, but the outcome $y^{s}_{i}$ then is sample from a Poisson distribution with mean $\mu({x}^{s}_{i},e\left({x}^{s}_{i};\widehat{\gamma}^{s}\right))$, which is the fitted value from a GEE model with $y^{s^{\prime}}_{i}$ with the inclusion of covariates $d^{s}_{i}$, the three-order spline of estimated GPS.
@@ 4.1 Non-causal empirical studies
There is a vast number of empirical studies that use non-causal statistical analysis methods to study the association of mobility with COVID-19 transmission as measured by either the number of COVID-19 cases or the virus reproduction number. Generally, most studies conclude that an increase in mobility is positively associated with an increase in COVID-19 transmission, and restrictions on mobility are associated with reductions in COVID-19 transmission. Studies using statistical correlation methods are prevalent in the literature. For example, Jia et al. (2020) reported a positive linear correlation between mobility and case incidence in China at the early stages of the pandemic. Badr et al. (2020) analysed mobility patterns in the USA and also reported positive linear correlations between mobility and COVID-19 case rates. Kissler et al. (2020) used Pearson’s linear correlation coefficient and reported that their estimates of COVID-19 prevalence was negatively associated with reductions in mobility in their analysis of New York City.
Studies adopting different forms of multivariate regression methods are also prevalent, and here we present a selection of examples. Li et al. (2021) used multivariate linear regression and reported that the reproduction number of the virus increased with higher levels of mobility at retail and recreation places, workplaces, and transit stations in the UK. Kraemer et al. (2020) applied a generalized linear regression model framework and found that mobility was positively correlated with case incidence in China prior to travel restrictions, but the relationship was no longer statistically significant after the implementation of restrictions. McGrail et al. (2020) undertook a cross-country analysis of 134 countries and 50 US states. The authors used a generalized linear mixed effects model and reported that mobility and infection rates reduced after the imposition of social distancing restrictions. Islam et al. (2020) also performed a cross-country analysis of 149 countries using a two-stage linear regression analysis. It was reported that case incidence reduced with the imposition of public transport closures, school closures, work closures, restrictions on mass gatherings, and blanket lockdowns. Xiong et al. (2020) reported a positive relationship between mobility and COVID-19 cases in the US using structural equation modelling with dynamic panel and time-varying effects at the level of US counties. Manzira et al. (2022) used a multivariate time series regression model with autoregressive lag effects and reported a positive association between mobility and the incidence of COVID-19 in the city centre areas of Dublin, Ireland.
@@ 4.2 Causal empirical studies
There are a limited number of studies that adopt causal analysis methods to quantify the impact of mobility on COVID-19 case transmission. Romero García et al. (2023) undertook the only known analysis of the impact of metro travel on COVID-19 case rates. Romero García et al. (2023) performed a small-scale analysis of the impact of the locations of subway stations and hospitals on COVID-19 cases in Valencia Spain. They first partitioned the region using Voroni diagrams using subway locations and hospitals as segmentation variables, and then applied the Granger causality test. A positive association was reported between interchange subway stations and COVID-19 cases, while transmission from hospital locations generated more mixed results.
Steiger et al. (2021), Nugent and Balzer (2023), Cho et al. (2022), and Chernozhukov et al. (2021) used aggregate indicators of mobility from mobile phone data records unattributed to specific travel modes to assess causality between mobility and COVID-19 cases. Steiger et al. (2021) estimated the causal impact of factors affecting the number of new cases of COVID-19 in Germany using a DAG and a negative binomial regression model. The authors found a positive causal impact between mobility in retail and recreational spaces and workplaces and COVID-19 cases, and a negative causal impact for essential shopping and mobility within residential areas. Other influential factors included weather, socio-demographic characteristics, and public awareness of COVID-19. Nugent and Balzer (2023) adopted the causal modified treatment policy method with the Super Learner machine learning algorithm to quantify the impact of mobility on COVID-19 cases in US counties. The authors found that unadjusted estimates suggested that decreasing mobility is causally linked to decreasing COVID-19 cases, but found that their adjusted estimates gave mixed results indicating potential confounding in the unadjusted estimates, which were most likely attributed to unknown differences in county characteristics. Cho et al. (2022) also analysed the causal relationship between COVID-19 cases and mobility in the US. They adopted an ensemble empirical mode decomposition method with causal decomposition and reported that mobility levels have a causal relationship with long-term variations in COVID-19 cases rather short-term variations. Chernozhukov et al. (2021) used a casual structural model framework to test a range of counterfactual experiments including mobility changes and their impact on COVID-19 cases in the US. The authors find that reductions in mobility due to stay at home orders and business closures were effective in reducing case numbers, while the impacts of reductions in mobility due to school closures were inconclusive.
@@ 4.3 Data
The Community of Metros (COMET) benchmarking group administered by the Transport Strategy Centre at Imperial College London have provided ridership data from 8 international member metro networks. In keeping with commercial confidentiality requirements, the metro identities have been anonymized and are referred to with a numerical index ranging from 1 to 8 throughout this paper. The ridership data report the numerical ridership figures for each day. The data are recorded from before the onset of the COVID-19 pandemic to early 2022, with the earliest and latest dates in the data set being 1 March 2020 and 23 January 2022.
@@
Ridership ($log$ scale) Number of passengers per day in log scale 10.17 15.15 13.02 1.07
@@ 4.4 Results
To estimate the dose-response relationship between the ridership and COVID-19 cases, we employed the proposed Bayesian method to analyze two different outcomes: the raw case counts, modeled using a log link with the Poisson mean-variance relationship, and the log-transformed case counts, $log(x+1)$, modeled under an identity with the Gaussian mean-variance relationship. For both outcomes, we specified an independent working correlation structure and included all relevant confounders along with a cubic spline function of the GPS as covariates. Similar to the simulation, for the GPS model, we used a GEE model that adjusts for all time-varying confounders. We generated a total of 5,000 posterior samples. For the DP method, we generate a synthetic dataset of size $J=200$ using a procedure similar to that employed in the simulation studies. Due to the relatively small sample size, the stick-breaking weights in the DP decreased rapidly, falling below $10^{-8}$ after approximately 200 iterations. Figure 4 shows the posterior predictive dose-response curves for WOR and COV methods. When using the COV method, we observe an increasing dose-response trend for both outcomes, suggesting that higher metro ridership is associated with a greater rise in COVID-19 cases. However, the variability in the estimates is higher in certain dose ranges (e.g, 10-12 and 14–15). The DP approach shows more stable curve as the synthetic dataset have a larger sample. When using case counts as the outcome, the COV method shows a generally exponential increase in cases as the logarithm of ridership rises. In contrast, the DP method appears to overcompensate in certain regions, resulting in some non-smooth curves likely because some portion of the resampled data is drawn from areas that deviate from the main trend. The WOR method, on the other hand, shows an initial rise in case counts that eventually plateaus and then slightly declines. In regions with sparse data, the DP approach produces more stable estimates, suggesting improved performance in low-density areas. We also report the estimated APOs at various doses of the posterior sample in Table 3. We choose the dose level at 11, 13 and 15, which corresponds to 2.5%, 50% and 97.5% quantiles of the logarithm of ridership respectively. Similar to the observation in Figure 4, the COV method reveals that the number of cases increases as the logarithm of ridership rises. On average, the number of cases increases from about 7330 to 59000 when dose level changes from 11 to 15, using $\log$ (case $+1$) as the outcome. When modeling the outcome as the count data, the number of cases increases by an even higher extent, from approximately 500000 to 5000000. This increase is attributed to the fact that exponential change when modeling the count data. However, the WOR method shows a different trend. While the number of cases initially increases, it begins to decline as the dose level rises from 13 to 15. In summary, the number of cases tends to increase as the dose level—representing public transport usage—rises, but this relationship seems not strictly linear. While both the log-transformed and count-based outcome suggest a strong upward trend in case numbers at lower dose regions, the log-transformed model shows a more moderate increase due to its smoothing effect, whereas the count model captures a steeper, exponential rise.
@@
Figure 4: Application: Posterior dose-response curves using WOR (right panel) and COV (left panel) approaches WOR and COV approaches, with outcomes defined as $\log$(case $+1$) and cases.
@@
COV-BB Posterior median $5.90\times 10^{5}$ $1.63\times 10^{6}$ $5.41\times 10^{6}$
@@
WOR-BB Posterior median $1.19\times 10^{6}$ $1.90\times 10^{6}$ $1.52\times 10^{6}$

### v2

@@ 3.1 Example 1
$D$ is a continuous treatment variable, and $X_{1}$, $X_{2}$ are time-varying confounders and $U$ is a time-invariant confounder for $D$ with a nonlinear relationship in $Y$. The goal of this causal setting is to estimate the APO, i.e., the outcome at a specific dose level. We consider estimating the dose-response relationship based on a GEE model for GPS with the inclusion of covariates $X_{1it}$, $X_{2it}$ only. For the COV method, we fit a GEE model for the outcome of $\log Y_{i}$ with the inclusion of covariates $D_{it}$, the three-order spline of estimated GPS. As for WOR method, we fit the outcome mean with the three-order spline of $D_{1it}$ with the GPS as the weight. In the DP method, the new data are generated from the old dataset or new data (unseen) from the base measure, $G_{0}$. We generate these new data through residuals arising from the outcome model. In this case, we first draw each pair of $\left\{y^{s^{\prime}}_{it},({x}^{s}_{it},d^{s}_{it})\right\}$, $t=1,\ldots,T_{i},i=1\ldots,N$, from the empirical distribution as the DP with $\alpha=0$, and then obtain the GPS, $\widehat{e}\left(\bar{x}^{s}_{ik}\right)$ from a mixed effect model with the inclusion of covariates $x^{s}_{1it}$, $x^{s}_{2it}$ to account for the effect of the unmeasured confounder (Graham et al., 2014), refitted to the newly sampled $\left\{({x}^{s}_{i},d^{s}_{i})\right\}$ data set. Then we simulate $y_{i}^{s}=\left\{y_{it}^{s}\right\}_{t=1}^{T}$ from a DP model with the conditional base measure $G_{0}\equiv\mathcal{N}\left(\mu({x}^{s}_{i},\widehat{e}\left(\bar{x}^{s}_{i}\right)),1\right)$, where $\mu({x}^{s}_{i},\widehat{e}\left(\bar{x}^{s}_{i}\right))$ is the fitted value from a GEE model with $\log y^{s^{\prime}}_{i}$ with the inclusion of covariates $d^{s}_{i}$, the three-order spline of estimated GPS, $\widehat{e}\left(\bar{x}^{s}_{i}\right)$. In the stick breaking construction, we truncate the sample size at $J=500$. The weight is less than $10^{-4}$ after 500 iterations and can be deemed as negligible.
@@ 3.2 Example 2
Similar to the first example, we fit a Poisson GEE model for the outcome of $Y_{i}$ with the inclusion of covariates $D_{1it}$, the three-order spline of estimated GPS for COV method, and as for WOR method, we fit the outcome mean with the three-order spline of $D_{1it}$ with the GPS as the weight. For the DP approach, we use the same steps to sample the data with size 500, but the outcome $y^{s}_{i}$ then is sample from a Poisson distribution with mean $g(\mathbb{E}[Y_{i}\mid D_{i},\bar{X}_{i},\widehat{e}_{i}(\bar{X}_{i})])$, which is the fitted value from a GEE model with $y^{s^{\prime}}_{i}$ with the inclusion of covariates $d^{s}_{i}$, the three-order spline of estimated GPS.
@@ 4.1 Non-causal empirical studies
There is a vast number of empirical studies that use non-causal statistical analysis methods to study the association of mobility with COVID-19 transmission as measured by either the number of COVID-19 cases or the virus reproduction number. Generally, most studies conclude that an increase in mobility is positively associated with an increase in COVID-19 transmission, and restrictions on mobility are associated with reductions in COVID-19 transmission. Studies using statistical correlation methods are prevalent in the literature. For example, Jia et al. (2020) reported a positive linear correlation between mobility and case incidence in China at the early stages of the pandemic. Badr et al. (2020) analysed mobility patterns in the USA and also reported positive linear correlations between mobility and COVID-19 case rates. Kissler et al. (2020) used Pearson’s linear correlation coefficient and reported that their estimates of COVID-19 prevalence were negatively associated with reductions in mobility in their analysis of New York City.
Studies adopting different forms of multivariate regression methods are also prevalent, and here we present a selection of examples. Li et al. (2021) used multivariate linear regression and reported that the reproduction number of the virus increased with higher levels of mobility at retail and recreation places, workplaces, and transit stations in the UK. Kraemer et al. (2020) applied a generalized linear regression model framework and found that mobility was positively correlated with case incidence in China prior to travel restrictions, but the relationship was no longer statistically significant after the implementation of restrictions. McGrail et al. (2020) undertook a cross-country analysis of 134 countries and 50 US states. The authors used a generalized linear mixed effects model and reported that mobility and infection rates reduced after the imposition of social distancing restrictions. Islam et al. (2020) also performed a cross-country analysis of 149 countries using a two-stage linear regression analysis. It was reported that case incidence reduced with the imposition of public transport closures, school closures, work closures, restrictions on mass gatherings, and blanket lockdowns. Xiong et al. (2020) reported a positive relationship between mobility and COVID-19 cases in the US using structural equation modeling with dynamic panel and time-varying effects at the level of US counties. Manzira et al. (2022) used a multivariate time series regression model with autoregressive lag effects and reported a positive association between mobility and the incidence of COVID-19 in the city centre areas of Dublin, Ireland.
@@ 4.2 Causal empirical studies
There are a limited number of studies that adopt causal analysis methods to quantify the impact of mobility on COVID-19 case transmission. Romero García et al. (2023) undertook the only known analysis of the impact of metro travel on COVID-19 case rates. Romero García et al. (2023) performed a small-scale analysis of the impact of the locations of subway stations and hospitals on COVID-19 cases in Valencia, Spain. They first partitioned the region using Voronoi diagrams using subway locations and hospitals as segmentation variables, and then applied the Granger causality test. A positive association was reported between interchange subway stations and COVID-19 cases, while transmission from hospital locations generated more mixed results.
Steiger et al. (2021), Nugent and Balzer (2023), Cho et al. (2022), and Chernozhukov et al. (2021) used aggregate indicators of mobility from mobile phone data records unattributed to specific travel modes to assess causality between mobility and COVID-19 cases. Steiger et al. (2021) estimated the causal impact of factors affecting the number of new cases of COVID-19 in Germany using a DAG and a negative binomial regression model. The authors found a positive causal impact between mobility in retail and recreational spaces and workplaces and COVID-19 cases, and a negative causal impact for essential shopping and mobility within residential areas. Other influential factors included weather, socio-demographic characteristics, and public awareness of COVID-19. Nugent and Balzer (2023) adopted the causal modified treatment policy method with the Super Learner machine learning algorithm to quantify the impact of mobility on COVID-19 cases in US counties. The authors found that unadjusted estimates suggested that decreasing mobility is causally linked to decreasing COVID-19 cases, but found that their adjusted estimates gave mixed results indicating potential confounding in the unadjusted estimates, which were most likely attributed to unknown differences in county characteristics. Cho et al. (2022) also analysed the causal relationship between COVID-19 cases and mobility in the US. They adopted an ensemble empirical mode decomposition method with causal decomposition and reported that mobility levels have a causal relationship with long-term variations in COVID-19 cases rather short-term variations. Chernozhukov et al. (2021) used a causal structural model framework to test a range of counterfactual experiments including mobility changes and their impact on COVID-19 cases in the US. The authors find that reductions in mobility due to stay at home orders and business closures were effective in reducing case numbers, while the impacts of reductions in mobility due to school closures were inconclusive.
@@ 4.3 Data
The Community of Metros (COMET) benchmarking group administered by the Transport Strategy Centre at Imperial College London have provided ridership data from 8 international member metro networks. In keeping with commercial confidentiality requirements, the metro identities have been anonymized and are referred to with a numerical index ranging from 1 to 8 throughout this paper. The ridership data reports the numerical ridership figures for each day. The data are recorded from before the onset of the COVID-19 pandemic to early 2022, with the earliest and latest dates in the data set being 1 March 2020 and 23 January 2022.
@@
Ridership (loglog scale) Number of passengers per day in log scale 10.17 15.15 13.02 1.07
@@ 4.4 Results
To estimate the dose-response relationship between the ridership and COVID-19 cases, we employed the proposed Bayesian method to analyze two different outcomes: the raw case counts, modeled using a log link with the Poisson mean-variance relationship, and the log-transformed case counts, $log(x+1)$, modeled under an identity link with the Gaussian mean-variance relationship. For both outcomes, we specified an independent working correlation structure and included all relevant confounders along with a cubic spline function of the GPS as covariates. Similar to the simulation, for the GPS model, we used a GEE model that adjusts for all time-varying confounders. We generated a total of 5,000 posterior samples. For the DP method, we generate a synthetic dataset of size $J=200$ using a procedure similar to that employed in the simulation studies. Due to the relatively small sample size, the stick-breaking weights in the DP decreased rapidly, falling below $10^{-8}$ after approximately 200 iterations. Figure 4 shows the posterior predictive dose-response curves for WOR and COV methods. When using the COV method, we observe an increasing dose-response trend for both outcomes, suggesting that higher metro ridership is associated with a greater rise in COVID-19 cases. However, the variability in the estimates is higher in certain dose ranges (e.g, 10-12 and 14–15). The DP approach shows a more stable curve as the synthetic dataset have a larger sample. When using case counts as the outcome, the COV method shows a generally exponential increase in cases as the logarithm of ridership rises. In contrast, the DP method appears to overcompensate in certain regions, resulting in some non-smooth curves likely because some portion of the resampled data is drawn from areas that deviate from the main trend. The WOR method, on the other hand, shows an initial rise in case counts that eventually plateaus and then slightly declines. In regions with sparse data, the DP approach produces more stable estimates, suggesting improved performance in low-density areas. We also report the estimated APOs at various doses of the posterior sample in Table 3. We choose the dose level at 11, 13 and 15, which corresponds to 2.5%, 50% and 97.5% quantiles of the logarithm of ridership, respectively. Similar to the observation in Figure 4, the COV method reveals that the number of cases increases as the logarithm of ridership rises. On average, the number of cases increases from about 7330 to 59000 when dose level changes from 11 to 15, using $\log$ (case +1+1) as the outcome. When modeling the outcome as the count data, the number of cases increases by an even higher extent, from approximately 500000 to 5000000. This increase is attributed to the fact that exponential change when modeling the count data. However, the WOR method shows a different trend. While the number of cases initially increases, it begins to decline as the dose level rises from 13 to 15. In summary, the number of cases tends to increase as the dose level—representing public transport usage—rises, but this relationship is not strictly linear. While both the log-transformed and count-based outcome suggest a strong upward trend in case numbers at lower dose regions, the log-transformed model shows a more moderate increase due to its smoothing effect, whereas the count model captures a steeper, exponential rise.
@@
Figure 4: Application: Posterior dose-response curves using WOR (right panel) and COV (left panel) approaches WOR and COV approaches, with outcomes defined as $\log$(case +1+1) and cases.
@@
COV-$B$ Posterior median $5.90\times 10^{5}$ $1.63\times 10^{6}$ $5.41\times 10^{6}$
@@
WOR-$B$ Posterior median $1.19\times 10^{6}$ $1.90\times 10^{6}$ $1.52\times 10^{6}$